# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ II.4 Mapping onto a spin problem
[…]
In the CB regime, where the spin accumulation is small, the dynamics of the spin are dominated by Eq. 20 as long as $\theta\neq 0,\pi$. These two extreme cases are interesting though since they imply a simple relaxation of the $z$ component of the spin to the steady without any evolution of the $x$ and $y$ components, see Eq. LABEL:spin_equations. In the absence of driving, the exchange field will generate a spin splitting of the QI, a Zeeman splitting, even at $B=0$. When the driving is turned on, both the rotation and accumulation term will exhibit time-dependent behavior that will yield FLT and STT resonance signals in the frequency domain respectively.
@@ III.1 Rabi rate
In previous studies, the Rabi frequency was introduced as a mathematical consequence of the model, conveying that the modulation of the tunneling barrier generates a driving of reasonable amplitude [27, 35, 36]. However, this analysis does not provide an intuitive understanding of the actual driving effect on the spin, thus it is useful to cast the driving mechanism into a more intuitive physical picture.
Here, we want to address the definition of the Rabi frequency by looking at the system of equations for the spin components for each of the two driving regimes considered in this work: FLT and STT. For the FLT-driven regime, which is presence in the CB regime, the simulations show the typical behavior of a spin under a time-dependent magnetic field, i.e. the spin performs Rabi oscillations. This points to the exchange field as the dominant source of driving. Contrary, in the STT-driven regime, presence above the thresholds that limit the CB regime, the spin only decays to a polarized steady state while performing Larmor oscillations in the plane perpendicular to the quantization axis of the spin. This is the result of the accumulation term being a dominant influence on the spin dynamics.
@@ III.1.1 Rabi rate in the FLT-driven regime
There are two essential ingredients for the Rabi rate to emerge naturally in the QME: the spin polarization of one of the leads and the driving amplitude. We already showed the consequence of introducing polarization on the rates by mapping the QME onto a spin problem in Eqs. LABEL:spin_equations: the spin accumulation and rotation terms emerge as convenient descriptors for the spin dynamics. In the CB regime, the spin accumulation is rather small and can be partly discarded. Thus, only the rotation and relaxation terms strongly influence the time evolution of the spin.
When the exchange field and the spin are forming an angle $\theta\neq 0,\pi$, the exchange field creates a torque-like force on the initial spin that, together with the spin relaxation, will evolve the QI to the steady state. If this FLT produced by the exchange field is time-dependent, it is easy to conclude that it will induce a Rabi oscillation during this relaxation process, $\pi/2$ radians dephased respect to the driving term [24]. The strength of this Rabi rate will be given by the component of the exchange field torquing the spin, which is:
$\Omega_{\mathrm{FLT}}\propto B_{\mathrm{exch}}\cos(\pi/2-\theta).$
Looking at Eq. 22 we can easily identify
@@
This simple equation of the Rabi rate shows its physical meaning clearly as the projection of the exchange field on the perpendicular axis to the spin quantization in the plane where the spin and the exchange field lie. A comparison between QME and the optical Bloch equations also provides the same identification of the Rabi rate as the imaginary part of the rate associated to a spin-flip since only the imaginary part can provide Rabi oscillations. As a clear consequence of this definition, the Rabi rate is maximized at $\pi/2+m\pi$ with $m$ being an integer. However, this does not directly translate into a higher ESR signal amplitude because of the homodyne detection mechanism: the ESR signal measured by the electric current is a measurement of the $z$ component of the spin, as we will show in the simulation results.
@@
Therefore, the exchange field provides zero-order and higher harmonic signals directly, which is in line with recent experiments on a strongly driven pentacene molecule spin-1/2 system [24, 69]. The $n$-th harmonic signal appears at a frequency of $f_{1\omega}/n$, with $f_{1\omega}$ being the first harmonic resonance frequency and $n>0$ the harmonic number. In this way, the second harmonic signal provides an additional resonance at $f_{1\omega}/2$ in the frequency spectrum. This can be easily verified since this 2th harmonic signal will emerge when $22\hbar\omega=4\pi\hbar f_{2\omega}=\Delta_{eg}\implies f_{2\omega}=\Delta_{eg}/2\hbar=f_{1\omega}/$ 2. If the frequency is fixed whilst sweeping the magnetic field, the $n$-th resonance magnetic field will roughly follow the relation $nB_{1\omega}$ with $B_{1\omega}$ the magnetic field of the first resonance [24, 69]. The zero-harmonic feature is a special case, and it is found at zero frequency in the frequency domain and near zero in the magnetic field domain.
In Ref. [24], the authors observed a peak near zero magnetic field with an intensity comparable to or slightly higher than that of the first harmonic peak. Notably, this zero-field peak appears even in the absence of driving. At first glance, based on the equations above, this result could be explained by the 0-th harmonic resonance, as both emerge at zero driving and require only a spin-polarized tip. However, the results in [24] were obtained in the STT-driven regime, making them more consistent with explanations based on spin bottlenecking [70] or the Hanle effect [43, 42]. Measurements in the FLT-driven regime of this phenomenon are still lacking, but they could offer coherent control of a single spin at zero AC driving, using only DC voltages.
Coming back to the equations, one can show that the resonance for the zero-harmonic signal occurs when $B=-B_{\mathrm{exch}}\cos\theta$ when the frequency is fixed. This suggests that this signal position can be used as a measurement of the exchange field. From our model’s perspective, the origin of this zero-field feature arises from spin-polarized electron precession in a magnetic field generated by the exchange interaction with the spin polarized tip. However, in the absence of driving, no ESR signal is expected for any harmonic when sweeping over the frequency for a fixed magnetic field, even at zero frequency, since the $\cos\omega t$ term vanishes in the rates.
@@ III.1.2 Rabi rate in the STT-driven regime
When increasing the DC bias $eV_{\mathrm{DC}}$ beyond the energy thresholds, $eV_{\mathrm{DC}}<\varepsilon$ and $eV_{\mathrm{DC}}>\varepsilon+U$, we are outside the CB regime and the electric current through the system drastically increases at constant tip-atom distance. In this regime, the role of the exchange field is suppressed by the dominance of the spin accumulation term and the rotation term starts to decrease logarithmically. This does not imply the lack of ESR signal when the QI is driven in the STT regime. In fact, we can define the Rabi rates in this regime since the rates in the spin accumulation term (Eq. 19) can be written as the real part of a rate linked to a Rabi process by applying the following relations:
@@
$=$ $\sin\theta\sum_{\alpha}P_{\alpha}\left[\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))\right]$
@@
$=$ ±$\pm\sum_{\alpha}\mbox{Re}(\Gamma_{vg,gv,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\sin\theta,$
@@
where the $+$ sign applies to $v=0$ while the -- sign is for $v=2$. The definitions of the Rabi rates above are easily confirmed by comparing our QME with the optical Bloch equation and considering that the Rabi rate is a complex number with a real and an imaginary part. If driving is applied, the Rabi rates above will modulate the spin accumulation, providing an ESR purely governed by the coherences [36]. On resonance, this forces the spin to undergo Larmor precession, in phase with the driving term, while it evolves to the steady state.
It is important to highlight that, in this regime, there is an inherent competition between the external magnetic field and the coupling to the spin polarized electrode, which determines where the spin will accumulate. If $B\gg\gamma_{L}^{0}$, the spin accumulation term is too weak to align the spin with the $z$-axis, and instead, the spin accumulates in the magnetic field orientation performing Larmor precession in the plane perpendicular to the spin’s quantization axis. Conversely, if $B\ll\gamma_{L}^{0}$, the spin accumulation term dominates, forcing the spin to align with the $z$-axis. This mechanism underlies the Hanle effect outside the CB regime [43, 42]. However, in both cases discussed, the sign of the spin accumulation determines whether the QI is spin polarized positively or negatively, disregarding the magnetic field orientation. In this work, we consider values of the coupling and magnetic field that faithfully reflect experimental conditions, by obtaining FLT Rabi frequencies in agreement with the experimental measurements, placing ourself in the $B\gg\gamma_{L}^{0}$ regime.
@@
A key consequence of Larmor precession occurring in the plane perpendicular to the spin’s quantization axis is that the populations remain unchanged even at resonance, a result that persists in the long-time limit. This implies the absence of Rabi oscillations since, essentially, the spin is being forcedly accumulated in the direction set by the external magnetic field. Moreover, the ESR signal does not follow the homodyne detection [34, 21], leading to a fully symmetric resonance [36]. Another key feature of this regime is that Rabi rates are proportional to the electric current, exhibiting anharmonic behavior, since both the current and the Rabi rate depend on the coupling to the tip. As a result, the STT-driven regime naturally gives rise to higher harmonic resonances [69].
III.1.3 Analytical form of the Rabi
We could encompass both the FLT and STT Rabi rates in a complex rate term as $\hbar\Omega(t)=\hbar(\Omega_{\mathrm{STT}}(t)+i\Omega_{\mathrm{FLT}}(t))=\sum_{v=0,2}\Gamma_{gv,ve}(t)$. Every rate is calculated numerically in the wide band limit, which mandates a cutoff in the integral. If the cutoff energy, $E_{c}$, is much larger than any other energy in the QI, i.e. $E_{c}\gg|\varepsilon|,|\varepsilon+U|$, the rates will converge to a value independent of the cutoff [38, 63, 42], which is the situation in all our calculations. In this case, at extremely low temperatures, $k_{B}T_{\alpha}\ll|\varepsilon-\mu_{\alpha}|,|\varepsilon+U-\mu_{\alpha}|$, we can assume that the Fermi distribution function behaves as a step function:
$f_{\alpha}^{+}(\epsilon)$$\approx$ $1-\Theta(\epsilon-\mu_{\alpha})$
$f_{\alpha}^{-}(\epsilon)$$\approx$ $\Theta(\epsilon-\mu_{\alpha}),$
with $\Theta(x)$ being the Heaviside step function. This approximation allows us to write the simplest [^5] An analytical expression for the imaginary part of the Rabi rate at finite temperatures can be expressed using digamma functions [71] but we are interested in the low temperature scenario. analytical form of the Rabi rate for a spin-$1/2$ system as:
@@ III.1.3 Analytical form of the Rabi
$+$ $\left.\frac{i}{2}\log\left(\frac{(\mu_{\alpha}-\varepsilon-U)^{2}+(\hbar/\tau_{c})^{2}}{(\mu_{\alpha}-\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right)\right].$
In the limit $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$, the real part of the Rabi frequency becomes close to zero in the CB regime, which results in nearly zero current as well. In contrast, outside of the CB regime, the sum of the $\arctan$ terms gives $\mbox{sign}(V_{\mathrm{DC}})\pi=\mbox{sign}(\mu_{L})\pi$, which depends on the sign of the chemical potential of the polarized electrode. On the other side, the imaginary part of Eq. 25 does not vanish in the CB regime, indicating its independence from the current, as we mentioned before.
@@ III.2 Spin relaxation time and quality factor $\Omega T_{2}$
Let us move on to the spin relaxation Eq. 18. Since our model does not consider any pure dephasing mechanism and we have not included cotunneling processes, the time independent part of $\tau_{\mathrm{rel}}$ is actually the coherence time $T_{2}$, which is purely lifetime-limited so that $T_{2}=2T_{1}$. Assuming low temperature again, one arrives at an analytical expression for $T_{2}$:
@@
An important feature of Eq. 26 is that it saturates rapidly as we increase $|V_{\mathrm{DC}}|$ outside of the CB regime to
@@
where the right chemical potential is grounded and its contribution can be discarded compared to the left one for $eV_{\mathrm{DC}}=\mu_{L}<\varepsilon$ and $eV_{\mathrm{DC}}>\varepsilon+U$. Equation 27 does not depend on the induced broadening $\hbar/\tau_{c}$, since the DC bias is above the energy thresholds and is essentially controlled by the time independent left coupling $\gamma^{0}_{L}$, which is where the driving is applied and the tip-atom distance is encoded.
In the CB regime, however, the coherence times depend on the induced broadening, and a finite value of $\hbar/\tau_{c}$ prevents $T_{2}$ from diverging. This feature is a clear limitation of the model, as it was mentioned before in the introduction, and going beyond it requires extending to higher orders in perturbation theory of the tunneling amplitude $t_{\alpha}$, which will lead to cotunneling [39, 27]. Future extensions of this model will aim to circumvent such limitation, nonetheless, we set $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$ so its influence on Eq. 26 is minimized. In this way, we can explore the behavior of $T_{2}$ in the CB regime although, at low DC voltage, we expect low current in line with previous studies [38, 62, 39, 63, 42] as the cotunneling regime becomes more significant in the deep CB regime.
The product $\Omega T_{2}$ per harmonic is the quality factor of a qubit, giving the number of Rabi oscillation the spin performs on resonance before reaching the steady state. We focus on the Rabi frequency related to $1\omega$, the first harmonic, under reasonable ESR conditions with $eV_{\mathrm{DC}}=\mu_{L}$ and the spin polarization and driving are only applied to the left lead. When the QI is driven under STT, the real part of the Rabi rate dominates, and we obtain:
@@
Equation 28 is proportional to the ESR amplitude and already suggests an optimal value for the driving. Additionally, it indicates that the ESR signal is independent of the couplings to the leads. If the driving is increased substantially, Eq. 28 will tend toward zero, and the ESR signal will diminish as well. Specifically, the driving value that maximizes the first harmonic signal is $\sqrt{2}$, yielding $\Omega_{\mathrm{STT}}^{1\omega}T_{2}=-\mathrm{sign}(V_{\mathrm{DC}})\sin\theta P_{L}\sqrt{2}/4$. Even in the ideal scenario of a fully polarized electrode and a magnetic field perpendicular to the exchange field, $|\Omega_{\mathrm{STT}}^{1\omega}T_{2}|=\sqrt{2}/4\approx 0.35$, meaning that in the STT-driven regime, we are unable to create a robust spin-qubit, as it will complete barely 1/3 of a Rabi oscillation before fully decaying to the steady state. This does not improve further if we consider the surviving Rabi rate of the FLT regime instead. Therefore, the spin is always overdamped in this scenario, and the ESR driven by STT is mainly useful for polarizing the QI as an initialization step, although the ESR signal can still be reasonably strong under suitable conditions [35, 36].
@@
$\Omega_{\mathrm{FLT}}^{1\omega}T_{2}\approx\frac{-i\gamma_{L}^{0}A_{L}P_{L}\sin\theta\log\left(\frac{(eV_{\mathrm{DC}}+\varepsilon+U)^{2}+(\hbar/\tau_{c})^{2}}{(eV_{\mathrm{DC}}+\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right)/2}{\gamma_{L}^{0}\left(1+\frac{A_{L}^{2}}{2}\right)\left[\pi-\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon}{\hbar/\tau_{c}}\right)+\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon-U}{\hbar/\tau_{c}}\right)\right]+\gamma_{R}^{0}\left[\pi+\arctan\left(\frac{\varepsilon}{\hbar/\tau_{c}}\right)-\arctan\left(\frac{\varepsilon+U}{\hbar/\tau_{c}}\right)\right]}.$ (29)
Equation 29 depends on both left and right couplings, unlike Eq. 27, since now $\varepsilon<eV_{\mathrm{DC}}<\varepsilon+U$ and both left and right terms are in a similar order of magnitude. It is also required to retain the broadening terms to avoid divergences as we are considering low temperatures. Increasing the coupling of the unpolarized electrode to values where $\gamma_{R}\gg\gamma_{L}$ suppresses the ESR signal since $\Omega_{\mathrm{FLT}}^{1\omega}T_{2}\rightarrow 0$, while increasing $\gamma_{L}$ causes the ESR amplitude to reach a saturation point that becomes independent of the couplings. The former implies influence from the substrate, indicating that extremely strong coupling to it leads to no ESR. Equation 29 depends on the DC voltage and enables further optimization of the qubit’s quality factor beyond the limit of 1/3 of the STT-driven regime, as we will detail in the results section. Both extreme situations between the couplings yield an electric current independent of the spin polarization, suggesting an ideal factor between $\gamma_{R}$ and $\gamma_{L}$ that maximizes the ESR signal in the FLT-driven regime.
@@
Outside of the CB regime, Eq. 23 does not fully vanish, but it logarithmically tends toward zero. However, the $T_{2}$ drops much faster, making Eq. 29 extremely small for any driven regime, preventing any quantum control. Thus, both FLT and STT regimes will give a measurable ESR signal in the electric current, though their origins and utilities differ significantly as we will show in the results section: STT-ESR provides a strong spin polarization of the QI, whereas for the FLT-ESR, the QI can be reliably used as a spin-qubit.

### v2

@@ II.4 Mapping onto a spin problem
where we have inserted the definition of the imaginary part of the rate. The name of this effective field originates from the fact that it arises due to virtual particle exchange between the electrode and the QI. This field is directly proportional to the electrode’s spin polarization and vanishes in the electron-hole symmetry point $eV_{\mathrm{DC}}=\varepsilon+U/2$, where we have assumed that the voltage is applied to the left electrochemical potential while the right one is grounded so $eV_{\mathrm{DC}}=\mu_{L}-\cancel{\mu_{R}}^{0}$. Furthermore, the exchange field vanishes in the absence of Coulomb interaction ($U=0$), demonstrating that its origin is a quantum many-body effect. It is important to notice that Eq. (22) does not change the charge of the QI, contrary to the spin accumulation term. Therefore, it can be finite even at zero voltage, indicating its independence from the DC electric current. However, Eq. (22) does depend on the charging energies $\varepsilon,\ \varepsilon+U$ and therefore can be understood as an exchange field due to virtual particle exchange with a spin-polarized lead.
In the CB regime, where the spin accumulation is small, the dynamics of the spin are dominated by Eq. (20) as long as $\theta\neq 0,\pi$. These two extreme cases are interesting though since they imply a simple relaxation of the $z$ component of the spin to the steady without any evolution of the $x$ and $y$ components, see Eqs. (17a). In the absence of driving, the exchange field will generate a spin splitting of the QI, a Zeeman splitting, even at $B=0$. When the driving is turned on, both the rotation and accumulation term will exhibit time-dependent behavior that will yield FLT and STT resonance signals in the frequency domain respectively.
@@ III.1 Rabi rate
In previous studies, the Rabi frequency was introduced as a mathematical consequence of the model, conveying that the modulation of the tunneling barrier generates a driving of reasonable amplitude [32, 37, 38]. While formally correct, that presentation does not give an immediate, physical intuition for how the tunneling-modulation acts on the spin. It is therefore instructive to recast the driving mechanism into a more intuitive picture and to examine explicitly the equations of motion for the spin components in the two driving regimes considered here: FLT and STT.
For the FLT-driven regime, which is predominant in the CB regime, the equations reproduce the typical behavior of a spin under a time-dependent magnetic field, the exchange field. Thus the impurity undergoes Rabi oscillations governed by the torque term (the cross product in Eq. (20)), and the observed ESR signal is strongly shaped by homodyne readout. In this regime the Rabi frequency is set by the component of the time-dependent exchange field that acts perpendicular to the instantaneous spin (i.e., the torque projection). By contrast, in the STT-driven regime, dominant above the energy thresholds that delimit the CB regime, the time-dependent spin accumulation governs the ESR signal. Physically, spin-transfer torque originates from the frequent repeated transfer of tunneling electrons with a specific polarization (orientation), generating a non-equilibrium, dissipative torque that drives the impurity incoherently toward a new polarized steady state. Since the populations are not being affected, the ESR signal is fully controlled by the coherences and proportional to $\sin\theta$; thus, it presents neither asymmetric contribution nor homodyne dependence.
@@ III.1.1 Rabi rate in the FLT-driven regime
There are two essential ingredients for the Rabi rate to emerge naturally in the QME: the spin polarization of one of the leads and the driving amplitude. We already showed the consequence of introducing polarization on the rates by mapping the QME onto a spin problem in Eqs. (17a) and (17b): the spin accumulation and rotation terms emerge as convenient descriptors for the spin dynamics. In the CB regime, the spin accumulation is rather small and can be partly discarded. Thus, only the rotation and relaxation terms strongly influence the time evolution of the spin.
When the exchange field and the spin form an angle $\theta\neq 0,\pi$, the former exerts a torque-like force on the initial spin, while spin relaxation ensures convergence to the steady state. If this FLT generated by the exchange field is time-dependent, it follows that Rabi oscillations will be induced during the relaxation process, $\pi/2$ radians out of phase with the driving term and the STT-driven mechanism [26, 73]. The corresponding Rabi rate is determined by the component of the exchange field acting as the torque on the spin, namely: $\Omega_{\mathrm{FLT}}\propto B_{\mathrm{exch}}\cos(\pi/2-\theta).$
Looking at Eq. (22) we can easily identify
@@
This simple expression for the Rabi rate highlights its physical meaning: it corresponds to the projection of the exchange field onto the axis perpendicular to the spin alignment (the eigenbasis axis) within the plane defined by the spin and the exchange field. A comparison between the QME and the optical Bloch equations leads to the same identification of the Rabi rate as the imaginary part of the spin-flip rate, since only the imaginary part generates oscillatory dynamics. As a direct consequence, the Rabi rate is maximized at $\pi/2+m\pi$, with $m$ an integer. However, this maximum does not necessarily translate into a stronger ESR signal amplitude because of the homodyne detection mechanism: the ESR signal measured through the tunneling current reflects only the $z$ component of the spin, as will be demonstrated in the simulation results.
@@
Therefore, the exchange field provides zero-order and higher harmonic signals directly, with the latter in agreement with recent experiments on a strongly driven pentacene molecule spin-1/2 system [77] The $n$-th harmonic signal appears at a frequency of $f_{1\omega}/n$, with $f_{1\omega}$ being the first harmonic resonance frequency and $n>0$ the harmonic number. In this way, the second harmonic signal provides an additional resonance at $f_{1\omega}/2$ in the frequency spectrum. This can be easily verified since this 2th harmonic signal will emerge when $2\hbar\omega=4\pi\hbar f_{2\omega}=\Delta_{eg}\implies f_{2\omega}=\Delta_{eg}/2\hbar=f_{1\omega}/2$. If the frequency is fixed whilst sweeping the magnetic field, the $n$-th resonance magnetic field will roughly follow the relation $nB_{1\omega}$ with $B_{1\omega}$ the magnetic field of the first resonance. The zero-harmonic feature is a special case, and it is found at zero frequency in the frequency domain and near zero in the magnetic field domain.
One can show that the resonance for the zero-harmonic signal occurs when $B=-B_{\mathrm{exch}}\cos\theta$ when the frequency is fixed. This suggests that this signal position can be used as a measurement of the exchange field. From our model’s perspective, the origin of this zero-field feature arises from spin-polarized electron precession in a magnetic field generated by the exchange interaction with the spin polarized tip. Nonetheless, experimental observation of this “resonance” is still lacking in the CB regime, but they could offer coherent control of a single spin at zero AC driving, using only DC voltages. In the absence of driving, no ESR signal is expected for any harmonic when sweeping over the frequency for a fixed magnetic field, even at zero frequency, since the $\cos(\omega t)$ term vanishes in the rates.
@@ III.1.2 Rabi rate in the STT-driven regime
When increasing the DC bias $eV_{\mathrm{DC}}$ beyond the energy thresholds, $eV_{\mathrm{DC}}<\varepsilon$ and $eV_{\mathrm{DC}}>\varepsilon+U$, we are outside the CB regime and the electric current through the system drastically increases at constant tip-atom distance. In this regime, the role of the exchange field is suppressed by the dominance of the spin accumulation term and the rotation term starts to decrease logarithmically. This does not imply the lack of an ESR signal when the QI is driven in the STT regime. In fact, we can define the Rabi rates in this regime since the rates in the spin accumulation term (Eq. (19)) can be written as the real part of a rate linked to a Rabi process by applying the following relations:
@@
$=$ $\sum_{\alpha}P_{\alpha}\left[\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))\right]\sin\theta$
@@
$=$ $\pm\sum_{\alpha}\mbox{Re}(\Gamma_{vg,gv,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\sin\theta,$
@@
where the ++ sign applies to $v=0$ while the - sign is for $v=2$. The definitions of the Rabi rates above are easily confirmed by comparing our QME with the optical Bloch equation and considering that the Rabi rate is a complex number with a real and an imaginary part. If driving is applied, the Rabi rates above will modulate the spin accumulation, providing an ESR purely governed by the coherences [38]. On resonance, this forces the spin to undergo Larmor precession, in phase with the driving term, while it rapidly evolves to the steady state. The precession occurs in the plane perpendicular to the eigenbasis axis, parallel to the magnetic field, so the populations, always defined with respect to that axis for us, remain unaltered even on resonance, a result that persists in the long time limit situation.
@@
As a consequence of the populations remaining unchanged even on resonance, the spin will not perform Rabi oscillations since, essentially, it is being forcedly locked (accumulated) in the direction set by the external magnetic field. Moreover, the ESR signal does not follow the homodyne detection [42, 22], leading to a fully symmetric resonance in the absence of exchange field [38]. Another key feature of this regime is that Rabi rates are proportional to the electric current, exhibiting anharmonic behavior, since both the current and the Rabi rate depend on the coupling to the tip. As a result, the STT-driven regime naturally gives rise to 0-th and higher harmonic resonances [77], similar to what the FLT-driven regime does. In Ref. [26], the authors observed a peak near zero magnetic field with an intensity comparable to or slightly higher than that of the first harmonic peak. Notably, this zero-field signal appears even in the absence of driving. Since these measurements were obtained in the STT-driven regime, explanations based on spin bottlenecking [78] or the Hanle effect [50, 46] are quite suitable. Physically, our 0-th harmonic is consistent with these interpretations.
It is important to emphasize the inherent competition between the external magnetic field and the coupling to the spin-polarized electrode, which determines where the spin will accumulate, as discussed in [41]. In the undriven regime, if $g\mu_{B}B\gg\gamma_{L}^{0}$, the spin-accumulation term is too weak to align the spin with the $z$-axis, and instead the spin accumulates approximately parallel to the magnetic field orientation. This alignment is not perfect: the residual influence of spin accumulation together with the torque exerted by the exchange field prevent the spin from pointing fully along $\mathbf{B}$. Conversely, if $g\mu_{B}B\ll\gamma_{L}^{0}$, the spin-accumulation term dominates, forcing the spin to align with the $z$-axis. This mechanism underlies the Hanle effect outside the CB regime [50, 46].
In both cases, the sign of the spin accumulation determines whether the QI is spin-polarized positively or negatively, independent of the magnetic-field orientation. In this work, we choose coupling and magnetic field values that realistically reflect experimental conditions, reproducing FLT Rabi frequencies consistent with experiment and placing us in the $g\mu_{B}B\gg\gamma_{L}^{0}$ regime. When AC driving is applied, however, the situation changes in the rotating frame and on resonance as $\theta$ varies from out-of-plane to in-plane orientations. In the latter case, the spin accumulates along the $z$-axis in the rotating frame, while for intermediate angles there is competition between accumulation along $z$ and parallel to the magnetic field.
III.1.3 Analytical form of the Rabi rate
[…]
@@ III.1.3 Analytical form of the Rabi rate
$+$ $\left.\frac{i}{2}\ln\left(\frac{(\mu_{\alpha}-\varepsilon-U)^{2}+(\hbar/\tau_{c})^{2}}{(\mu_{\alpha}-\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right)\right].$
In the limit $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$, the real part of the Rabi frequency becomes close to zero in the CB regime, which results in nearly zero current as well. In contrast, outside of the CB regime, the sum of the $\arctan$ terms gives $\pi\mbox{sign}(\mu_{L})=\pi\mbox{sign}(V_{\mathrm{DC}})$, which depends on the sign of the electrochemical potential of the polarized electrode. On the other side, the imaginary part of Eq. (25) does not vanish in the CB regime, indicating its independence from the current, as we mentioned before.
@@ III.2 Spin relaxation time and quality factor $\Omega T_{2}$
Let us move on to the spin relaxation Eq. (18). Since our model does not consider any pure dephasing mechanism and we have not included cotunneling processes, the time independent part of $\tau_{\mathrm{rel}}$ is actually the coherence time $T_{2}$, which is purely lifetime-limited so that $T_{2}=2T_{1}$. Assuming low temperature again, one arrives at an analytical expression for $T_{2}$:
@@
An important feature of Eq. (26) is that it saturates rapidly as we increase $|V_{\mathrm{DC}}|$ outside of the CB regime to
@@
where the right electrochemical potential is grounded and its contribution can be discarded compared to the left one for $eV_{\mathrm{DC}}=\mu_{L}<\varepsilon$ and $eV_{\mathrm{DC}}>\varepsilon+U$. Equation (27) does not depend on the induced broadening $\hbar/\tau_{c}$, since the DC bias is above the energy thresholds and is essentially controlled by the time independent left coupling $\gamma^{0}_{L}$, which is where the driving is applied and the tip-atom distance is encoded.
In the CB regime, however, the coherence times depend on the induced broadening, and a finite value of $\hbar/\tau_{c}$ prevents $T_{2}$ from diverging. This feature is a clear limitation of the model, as it was mentioned before in the introduction, and going beyond it requires extending to higher orders in perturbation theory of the tunneling amplitude $t_{\alpha}$, which will lead to cotunneling [45, 32]. Future extensions of this model will aim to circumvent such limitation, nonetheless, we set $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$ so its influence on Eq. (26) is minimized. In this way, we can explore the behavior of $T_{2}$ in the CB regime, although, at low DC voltage, we expect low current in line with previous studies [41, 70, 45, 71, 46, 47] as the cotunneling regime becomes more significant in the deep CB regime.
The product $\Omega T_{2}$ per harmonic is the quality factor of a qubit, giving the number of Rabi oscillation the spin performs on resonance before reaching the steady state. We focus on the Rabi frequency related to $1\omega$, the first harmonic, with $eV_{\mathrm{DC}}=\mu_{L}$, and the spin polarization and driving only applied to the left lead. When the QI is driven under STT, the real part of the Rabi rate dominates, and we obtain:
@@
Equation (28) is proportional to the ESR amplitude and already suggests an optimal value for the driving. Additionally, it indicates that the ESR signal is independent of the couplings to the leads. If the driving is increased substantially, Eq. (28) will tend toward zero, and the ESR signal will diminish as well. Specifically, the driving value that maximizes the first harmonic signal is $\sqrt{2}$, yielding $\Omega_{\mathrm{STT}}^{1\omega}T_{2}=-\mathrm{sign}(V_{\mathrm{DC}})\sin\theta P_{L}\sqrt{2}/4$. Even in the ideal scenario of a fully polarized electrode and a magnetic field perpendicular to the exchange field, $|\Omega_{\mathrm{STT}}^{1\omega}T_{2}|=\sqrt{2}/4\approx 0.35$, meaning that in the STT-driven regime, we are unable to create a robust spin-qubit, as it will complete barely 1/3 of a Rabi oscillation before fully decaying to the steady state. This does not improve further if we consider the surviving Rabi rate of the FLT regime instead. Therefore, the spin is always overdamped in this scenario, and the ESR driven by STT is mainly useful for polarizing the QI as an initialization step, although the ESR signal can still be reasonably strong under suitable conditions [37, 38].
@@
$\Omega_{\mathrm{FLT}}^{1\omega}T_{2}\approx\frac{-iA_{L}\gamma_{L}^{0}P_{L}\sin\theta\ln\left(\frac{(eV_{\mathrm{DC}}+\varepsilon+U)^{2}+(\hbar/\tau_{c})^{2}}{(eV_{\mathrm{DC}}+\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right)/2}{\gamma_{L}^{0}\left(1+\frac{A_{L}^{2}}{2}\right)\left[\pi-\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon}{\hbar/\tau_{c}}\right)+\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon-U}{\hbar/\tau_{c}}\right)\right]+\gamma_{R}^{0}\left[\pi+\arctan\left(\frac{\varepsilon}{\hbar/\tau_{c}}\right)-\arctan\left(\frac{\varepsilon+U}{\hbar/\tau_{c}}\right)\right]}.$
Equation (29) depends on both left and right couplings, unlike Eq. (27), since now $\varepsilon<eV_{\mathrm{DC}}<\varepsilon+U$ and the terms left and right are both of a similar order of magnitude. It is also required to retain the broadening terms to avoid divergences as we are considering low temperatures. Increasing the coupling of the unpolarized electrode to values where $\gamma_{R}^{0}\gg\gamma_{L}^{0}$ suppresses the ESR signal since $\Omega_{\mathrm{FLT}}^{1\omega}T_{2}\rightarrow 0$, while increasing $\gamma_{L}$ causes the ESR amplitude to reach a saturation point that becomes independent of the couplings. The former implies influence from the substrate, indicating that extremely strong coupling to it leads to no ESR. Equation (29) depends on the DC voltage and enables further optimization of the qubit’s quality factor beyond the limit of 1/3 of the STT-driven regime, as we will detail in the results section. Both extreme situations between the couplings yield an electric current independent of the spin polarization, suggesting an ideal factor between $\gamma_{R}^{0}$ and $\gamma_{L}^{0}$ that maximizes the ESR signal in the FLT-driven regime.
@@
Outside of the CB regime, Eq. (23) does not fully vanish, but it logarithmically tends toward zero. However, the $T_{2}$ drops much faster, making Eq. (29) extremely small for any driven regime, preventing any quantum control. Thus, both FLT and STT regimes will give a measurable ESR signal in the electric current, though their origins and utilities differ significantly, as we will show in the results section: STT-ESR provides a strong spin polarization of the QI, whereas for the FLT-ESR, the QI can be reliably used as a spin-qubit.